b1az3:

Which of the following equations is of a parabola with a vertex at (0, -5)? y = x2 - 5 y = x2 + 5 y = ( x + 5) 2 y = ( x - 5) 2

AZ:

y = (x-h)^2 + k where (h, k) is the vertex

AZ:

so they told you that your vertex is (0, -5) replace h with 0 and replace k with -5 what do you get?

b1az3:

y = ( x - 5) 2

AZ:

h is 0 k is -5 y = (x - h)^2 + k

AZ:

Do you need colors? \(\large\color{red}{h}\) is \(\color{red}{0}\) \(\large\color{orange}{k}\) is \(\color{orange}{-5}\) \( y = (x - \color{red}{h})^2 + \color{orange}{k}\)

b1az3:

x^2-5

AZ:

That's much better, Kyle!

b1az3:

thanks bro

AZ:

No problem! Hopefully the colors helped you see where what should go where